OpenStudy (anonymous):

Let f(x)=-4x+7 and g(x)=10x-6.find f(g(x))

OpenStudy (zzr0ck3r):

\(f(x) = -4(x)+7\\\color{blue}{g(x)= 10x-6\\}\) \(f(g(x)) = f(\color{blue}{g(x)})=-4(\color{blue}{g(x)})+7=-4(\color{blue}{10x-6})+7\) Simplify.